# From interpretability to inference: an estimation framework for universal approximators

### v5

@@ Abstract
We present a novel framework for estimation and inference for the broad class of universal approximators. Estimation is based on the decomposition of model predictions into Shapley values. Inference relies on analyzing the bias and variance properties of individual Shapley components. We show that Shapley value estimation is asymptotically unbiased, and we introduce Shapley regressions as a tool to uncover the true data generating process from noisy data alone. The well-known case of the linear regression is the special case in our framework if the model is linear in parameters. We present theoretical, numerical, and empirical results for the estimation of heterogeneous treatment effects as our guiding example.
@@ 2 Related literature
Regarding interpretability, a primary concern with the use of machine learning models is the black box critique. That is, given an optimized, or trained, model, its input-output relations are not directly accessible even if one has full access to the model. This is because the internal parameter space of a model generally is degenerate removing any intrinsic meaning of individual parameters (see next section for details). The attribution of importance scores to the individual predictors entering a model is a general approach to address this problem (see Molnar (2019) for an overview). These scores can be global or local. Global scores assign a single value to each variable across the input domain. A prominent example is variable importance for tree-based models (Friedman et al., 2009; Kazemitabar et al., 2017) Local measures provide scores for individual predictions (e.g. LIME (Ribeiro et al., 2016) or Shapley values (Strumbelj and Kononenko, 2010)). Local measures can always be aggregated to give a global measure, hence, they carry more information content. Lundberg and Lee (2017) show that Shapley values unify several local explanation measures. Together with the appealing analytical properties stemming from their game theoretic origin, it can be argued that Shapley values are the preferred measure to assess the importance of variables in a wide range of supervised learning settings. However, Sundararajan and Najmi (2019) showed that, despite their axiomatic definition, the operationalization of Shapley values often can render uniqueness results meaningless, and can lead to counter-intuitive attributions. As a minor contribution, we will slightly extend their results by providing an additional criterion between models, which is usually fulfilled in empirical settings, to alleviate their concerns.
@@ 3.1 Modelling setting and notation
$y_{i}\,=\,\hat{f}\big{(}x_{i};\theta)\,+\,\epsilon_{i}\,,$ (1)
with the mean-zero residual vector $\epsilon_{i}\in\mathbb{R}^{r}$. We only consider the one-dimensional case $r=1$ without loss of generality. The optimization problem for our models is to minimize an empirical risk $R^{e}=\frac{1}{m}\sum_{i=1}^{m}||\hat{y_{i}}-y_{i}||=\frac{1}{m}\sum_{i=1}^{m}||\epsilon_{i}||$, where |||||$\cdot$||| | ⋅ | | is a distance measure depending on the model. This describes a regression setting, while most aspects discussed here can be straightforwardly transferred to classification problems by varying the risk function (Vapnik (1999)).
@@
We use the index convention that $i,j\in\{1,\dots,m\}$ refer to individual observations (rows of $x$) and $k,l\in\{1,\dots,n\}$ to variable dimensions (columns of $x$). No index refers to the whole dataset. Estimated quantities are hatted, except decompositions or their components $\phi/\Phi$ for simplicity. Primed inputs, e.g. $x^{\prime}$, refer to variable sets out of the set of possible variable coalitions $\mathcal{C}(x)$, where each variable is allowed to enter at most once. We mostly refrain from using set braces for a simplified notation, and $|x^{\prime}|\leq n$ is the number of variables in $x^{\prime}$ (also indexed by $k,l$ if needed). Super-scripts $S$ refer to ‘Shapley-related’ quantities, which will be clear from the context. Star-quantities ($\phi^{\star}/\Phi^{\star}$) refer to (unobserved) true values.
@@ 3.2 The key assumption of learning
The main assumption which many of our results are based on is that statistical learning is possible. That is, the empirical risk $R^{e}$ converges uniformly in probability to the minimally achievable loss given by the irreducible error $\eta$ with a rate $\xi_{ml}$, i.e. $R^{e}\sim m^{-\xi_{ml}}$. Concretely, the expected value of model predictions convergences uniformly in probability to the true DGP
$\lim_{m_{train}\to\infty}\mathbb{E}_{test}\big{[}||f-\hat{f}||\big{]}=0\;,\;\forall x_{i}\in\Omega_{test}\,.$ (2)
The expectation $\mathbb{E}_{test}$ refers to model predictions at points $x_{i}\in\Omega_{test}$ from a hold-out test sample after training on an independent training sample $\Omega_{train}$. Property (2) is called error consistency, and the model $\hat{f}$ a universal approximator if it holds. This will be our main prerequisite:
@@
We need the train-test split of the data, such that $\mathbb{E}[\epsilon_{i}]=0$ holds. This can be seen as follows. Assume that a training observation $i$ influences its own target prediction $\hat{f}(x_{i})$ by some value $|\delta_{i}|>0$ through the optimization process. That is, including $i$ in the training data moves its predicted value by $\delta_{i}$. Then, we have an unknown relationship $\epsilon_{i}(\delta_{i})$ with $\mathbb{E}[\epsilon_{i}(\delta_{i})]\neq 0$ in general, i.e. there is a wedge created by in-sample evaluation. Thus, all expectations will be taken over test samples. We will drop the $train$- and $test$-subscripts in most instances unless explicitly needed for clarity.
@@ 3.3 Model training and testing
An important question is what $K$ should be. Some of our main results will depend on $\xi_{ml}\geq\frac{1}{2}$. However, we often have $\xi_{ml}<\frac{1}{2}$ in nonparametric learning settings (Stone (1982)). The solution is to choose $K$ large enough to increase the speed of nonparametric learning using the training partitions relative to the parametric rate of $\xi_{p}=\frac{1}{2}$ on the test partitions. In particular, we can set
$K\,\geq\,\Big{\lceil}m^{1-2\xi_{ml}}\Big{\rceil}+1\,\equiv\,\overline{K}\,,$ (3)
where equality leads to the parametric rate for quantities evaluated on the test sets if $m^{1-2\xi_{ml}}$ is an integer. To provide some intuition for the above expression, we can set $\xi_{ml}=\xi_{p}=\frac{1}{2}$. This leads to $\overline{K}=2$, which is the smallest possible number of folds, splitting the data into one training and one test partition. Now, defining $\underline{K}\equiv\lfloor m^{1-2\xi_{ml}}\rfloor+1$, we arrive at
@@
(super-convergence) Let $\xi_{ml}\leq\frac{1}{2}$ be the convergence rate for training $\hat{f}$. If $K>\underline{K}$, the variance of an estimator $E(\hat{f},x_{i})$ linear in $\hat{f}$ coming from sample splitting vanishes asymptotically relative to that of a classical $\sqrt{m}$-estimator, e.g. that of a linear regression coefficient. The proof is given in the Appendix.
@@
Assumption 2: The training-test setting is always such that the model $\hat{f}$ converges at least with the parametric rate, i.e. $K\geq\overline{K}\Rightarrow\xi_{ml}^{test}\geq\xi_{p}=\frac{1}{2}$.
@@ 3.4 Model Shapley values
Equations 4 & 5 do not account for the case when variables jointly contribute to model predictions, i.e. when they are dependent or interact. This situation can be addressed by using the more general Shapley-Taylor decomposition proposed by Agarwal et al. (2019): the discrete set derivative of model $\hat{f}$ at point $x_{i}$ with respect to the set $x^{\prime}$ conditioned on $x^{\prime\prime}$ is defined as
@@
$\phi_{h}\big{(}\hat{f},x_{i}\,|\,x^{\prime}\big{)}\,=\,\begin{cases}\delta_{x^{\prime}}\hat{f}(x_{i}\,|\,\emptyset)&\quad\text{if}\quad|x^{\prime}|<h\,,\\ \frac{h}{n}\sum_{x^{\prime\prime}\subseteq\mathcal{C}(x)\setminus x^{\prime}}\frac{\delta_{x^{\prime}}\hat{f}(x_{i}\,|\,x^{\prime\prime})}{\binom{n-1}{|x^{\prime\prime}|}}&\quad\text{if}\quad|x^{\prime}|=h\,.\end{cases}$ (7)
@@
with $\phi_{0}$ being the same intercept as in (4). The second sum generalizes summation over variables to that over Shapley-Taylor components. We suppress the expansion order $h$ in what follows if not explicitly needed. Statistical models usually do not allow for missing inputs, i.e. $|x^{\prime}|<n$. We address this by integrating out variables excluded from the model over a background $x_{bg}$. The intercept $\phi_{0}$ is then the expected predicted value over $x_{bg}$, i.e. $\phi_{0}=\mathbb{E}[\hat{f}(\emptyset)]$. This provides the reference value against which each Shapley component is measured. Thus, the choice of the background data is important for the interpretation of Shapley components, which we will discuss below. [^3] The formal properties of Shapley values, our computational approach, and a comparison of the Shapley-Taylor decomposition of model predictions and the the Taylor expansion of a analytical function are given in the Online Appendix.
@@ 4.1 Shapley estimation
$\bar{\phi}_{x^{\prime}}^{R}\,=\,\mathbb{E}_{\Omega}\big{[}\phi_{i,x^{\prime}}^{R}\big{]}=\frac{1}{m}\sum_{i=1}^{m}\,\bar{\phi}_{i,x^{\prime}}^{R}\,.$ (10)
@@
(analytical continuity I) The Shapley decomposition $\Phi$ of a model $\hat{f}$ linear in parameters $\theta$, $\hat{f}(x)=x\theta$, is the model itself with $\Phi=\Phi_{1}$ (Eq. 4), and $\phi_{k}^{lin}=(x_{k}-\bar{x}_{k})\hat{\theta}_{k}$ (Corollary 1 in Lundberg and Lee (2017)). The proof is given in the Appendix.
@@ 4.1.1 Application: Shapley estimation for heterogeneous treatment effects
$=$ $\bigg{[}\sum_{k}\phi_{i,k}^{t=1}+\sum_{k,l;0>j>k}\phi_{i,kl}^{t=1}\bigg{]}-\bigg{[}\sum_{k}\phi_{i,k}^{t=0}+\sum_{k,l;0>j>k}\phi_{i,kl}^{t=0}\bigg{]}$ (12)
@@
$=$ $\phi_{i,t}^{t=1}+\sum_{k;k\neq t}\phi_{i,tk}^{t=1}\,=\,\phi_{t}+\sum_{k;k\neq t}\phi_{i,t*k}\,.\vspace*{-0.3cm}$ (14)
The second row inserts the definition of the Shapley-Taylor decomposition where single and double indexed terms correspond to the main effects ($|x^{\prime}|=1$) and pairwise interactions ($|x^{\prime}|=2$), respectively. Any term not involving the treatment $t$ cancels out in the third row, including the intercept $\phi_{0}$.
Going to the forth row, we set the background data against which variable coalitions are evaluated to $x_{bg}=(t=0,z)\equiv x_{bg}^{0}$, i.e. the control group, or a representative subsample or summary of the untreated population. This has two consequences for for the treatment function. First, all terms in the second bracket of (13) vanish when going to (14) due to the missingness property of Shapley values. Second, the first terms $\phi_{t}$ in (14), which is constant across subjects, is an estimate of the average treatment effect (ATE).
@@ 4.2 Shapley estimation uncertainty
     $\cdot$ split $x$ into $x_{train}^{s}$ and $x_{test}^{s}$ using cross-fitting
@@
        $\cdot$ draw bootstrap sample $x_{b}^{s}$ from $x_{train}^{s}$
@@
        $\cdot$ perform Shapley decomposition $\Phi^{b}(x_{test}^{s};\hat{f}_{b})$
@@
$\sqrt{m}\,\bar{\phi}_{x^{\prime}}\;\;\underset{\mbox{\scriptsize$m\rightarrow\infty$}}{\text{\Large$\longrightarrow$}}\;\;\mathcal{N}\big{(}\bar{\phi}_{x^{\prime}}^{\star},\sigma_{x^{\prime}}^{\star}\big{)}$ (16)
@@
(Sample split confidence intervals) Let $[\phi^{B}_{low},\phi^{B}_{high}]$ and $[\phi^{R}_{low},\phi^{R}_{high}]$ be confidence intervals at some level $\gamma$ from independent estimations, such as those resulting from bootstrap estimation (Algorithm 1) and sample split variation, respectively. Both intervals are median-centered at zero without loss of generality, then the joint confidence interval at level $\gamma$ is bounded as
$[\phi^{joint}_{low},\phi^{joint}_{high}]\;\subseteq\;[\phi^{B}_{low}+\phi^{R}_{low},\phi^{B}_{high}+\phi^{R}_{high}]\,.$ (17)
@@
The right-hand side of Equation 17 provides a conservative bound assuming that large variations in the estimation of $\phi^{B}$ coincides with large variations of $\phi^{R}$. However, when equating these quantities with sampling and sample split estimates, respectively, we often observe that the latter has considerably smaller variation than the former, i.e. $\phi^{R}_{high}-\phi^{R}_{low}\ll\phi^{B}_{high}-\phi^{B}_{low}$, such that right-hand side of (17) provides a practical adjustment. We will see in the empirical case study that such adjustments accounting for split uncertainty are mostly small, especially for aggregate measures where variation from the observation level cancels out.

### v6

@@ Abstract
We present a novel framework for estimation and inference with the broad class of universal approximators. Estimation is based on the decomposition of model predictions into Shapley values. Inference relies on analyzing the bias and variance properties of individual Shapley components. We show that Shapley value estimation is asymptotically unbiased, and we introduce Shapley regressions as a tool to uncover the true data generating process from noisy data alone. The well-known case of the linear regression is the special case in our framework if the model is linear in parameters. We present theoretical, numerical, and empirical results for the estimation of heterogeneous treatment effects as our guiding example.
@@ 2 Related literature
Regarding interpretability, a primary concern with the use of machine learning models is the black box critique. That is, given an optimized, or trained, model, its input-output relations are not directly accessible even if one has full access to the model. This is because the internal parameter space of a model generally is degenerate removing any intrinsic meaning of individual parameters (see next section for details). The attribution of importance scores to the individual predictors entering a model is a general approach to address this problem (see Molnar (2019) for an overview). These scores can be global or local. Global scores assign a single value to each variable across the input domain. A prominent example is variable importance for tree-based models (Friedman et al., 2009; Kazemitabar et al., 2017) Local measures provide scores for individual predictions (e.g. LIME (Ribeiro et al., 2016) or Shapley values (Strumbelj and Kononenko, 2010)). Local measures can always be aggregated to give a global measure, hence, they carry more information content. Lundberg and Lee (2017) show that Shapley values unify several local explanation measures. Together with the appealing analytical properties stemming from their game theoretic origin, it can be argued that Shapley values are the preferred measure to assess the importance of variables in a wide range of supervised learning settings. However, Sundararajan and Najmi (2019) showed that, despite their axiomatic definition, the operationalization of Shapley values often can render uniqueness results meaningless, and can lead to counter-intuitive attributions. We provide a condition, which usually is fulfilled in empirical settings, to alleviate their concerns. [^2] This supplementary result is stated in the Online Appendix where we discuss the properties and computation of model Shapley values which are not the focus of our main study.
@@ 3.1 Modelling setting and notation
$y_{i}\,=\,\hat{f}\big{(}x_{i};\theta)\,+\,\epsilon_{i}\,,$
with the mean-zero residual vector $\epsilon_{i}\in\mathbb{R}^{r}$. We only consider the one-dimensional case $r=1$ without loss of generality. The optimization problem for our models is to minimize an empirical risk $R^{e}=\frac{1}{m}\sum_{i=1}^{m}||\hat{y_{i}}-y_{i}||=\frac{1}{m}\sum_{i=1}^{m}||\epsilon_{i}||$, where |||||$\cdot$||| | ⋅ | | is a distance measure depending on the model. This describes a regression setting, while most aspects discussed here can be straightforwardly transferred to classification problems by varying the risk function (Vapnik, 1999).
@@
We use the index convention that $i,j\in\{1,\dots,m\}$ refer to individual observations (rows of $x$) and $k,l\in\{1,\dots,n\}$ to variable dimensions (columns of $x$). No index refers to the whole dataset. Sample averages are barred. Estimated quantities are hatted, except decompositions or their components $\phi/\Phi$ for simplicity. Primed inputs, e.g. $x^{\prime}$, refer to variable sets out of the set of possible variable coalitions $\mathcal{C}(x)$, where each variable is allowed to enter at most once. We mostly refrain from using set braces for a simplified notation, and $|x^{\prime}|\leq n$ is the number of variables in $x^{\prime}$ (also indexed by $k,l$ if needed). Super-scripts $S$ refer to ‘Shapley-related’ quantities, which will be clear from the context. Star-quantities ($\phi^{\star}/\Phi^{\star}$) refer to (unobserved) true values.
@@ 3.2 The key assumption of learning
The main assumption which many of our results are based on is that statistical learning is possible. That is, the empirical risk $R^{e}$ converges uniformly in probability to the minimally achievable loss given by the irreducible error $\eta$ with a rate $\xi_{\mathrm{ml}}$, i.e. $R^{e}\sim m^{-\xi_{\mathrm{ml}}}$. Concretely, the expected value of model predictions convergences uniformly in probability to the true DGP
$\lim_{m_{\mathrm{train}}\to\infty}\mathbb{E}_{\mathrm{train}}\big{[}||f(x_{i})-\hat{f(}x_{i})||\big{]}=0\;,\;\forall x_{i}\in x_{\mathrm{test}}\,.$ (1)
The expectation $\mathbb{E}_{\mathrm{train}}$ refers to any independent training data $x_{\mathrm{train}}\subset\Omega$ of size $m_{\mathrm{train}}$ (training sample), while model evaluation is based on an separate hold-out sample $x_{\mathrm{test}}\subset\Omega$ (test sample). Property (1) is called error consistency, because the expected error $\epsilon$ vanishes everywhere despite the presence of the irreducible noise $\eta$. The model $\hat{f}$ is a universal approximator if (1) holds. [^3] Many popular machine learning models, like different types of ANN, RF, and SVM have been shown to be universal approximators. See for example Cybenko (1989); Geman et al. (1992); Farago and Lugosi (1993); Steinwart (2002); Steinwart and Scovel (2007); Christmann and Steinwart (2008); Biau (2012); Scornet et al. (2014); Andoni et al. (2014) and references therein. This will be our main prerequisite:
@@
We need the train-test split of the data, such that $\mathbb{E}_{\mathrm{train}}[\epsilon_{i}]=0,\;\forall x_{i}\in x_{\mathrm{test}}$. This can be seen as follows. Assume that a training observation $i$ influences its own target prediction $\hat{f}(x_{i})$ by some value $|\delta_{i}|>0$ through the optimization process. That is, including $i$ in the training data moves its predicted value by $\delta_{i}$. Then, we have an unknown relationship $\epsilon_{i}(\delta_{i})$ with $\mathbb{E}_{\mathrm{train}}[\epsilon_{i}(\delta_{i})]\neq 0$ in general, i.e. there is a wedge created by in-sample evaluation. Thus, all model expectations will be taken as in (1), and we will drop the $\mathrm{train}$- and $\mathrm{test}$-subscripts in most instances unless explicitly needed for clarity.
@@ 3.3 Model training and testing
An important question is what $K$ should be. Some of our main results will depend on $\xi_{\mathrm{ml}}\geq\frac{1}{2}$. However, we often have $\xi_{\mathrm{ml}}<\frac{1}{2}$ in nonparametric learning settings (Stone (1982)). The solution is to choose $K$ large enough to increase the speed of nonparametric learning using the training partitions relative to the parametric rate of $\xi_{p}=\frac{1}{2}$ on the test partitions. In particular, we can set
$K\,\geq\,\Big{\lceil}m^{1-2\xi_{\mathrm{ml}}}\Big{\rceil}+1\,\equiv\,\overline{K}\,,$ (2)
where equality leads to the parametric rate for quantities evaluated on the test sets if $m^{1-2\xi_{\mathrm{ml}}}$ is an integer. To provide some intuition for the above expression, we can set $\xi_{\mathrm{ml}}=\xi_{p}=\frac{1}{2}$. This leads to $\overline{K}=2$, which is the smallest possible number of folds, splitting the data into one training and one test partition. Now, defining $\underline{K}\equiv\lfloor m^{1-2\xi_{\mathrm{ml}}}\rfloor+1$, we arrive at
@@
(super-convergence) Let $\xi_{\mathrm{ml}}\leq\frac{1}{2}$ be the convergence rate for training $\hat{f}$. If $K>\underline{K}$, the variance of an estimator $E(\hat{f},x_{i})$ linear in $\hat{f}$ coming from sample splitting vanishes asymptotically relative to that of a classical $\sqrt{m}$-estimator, e.g. that of a linear regression coefficient. The proof is given in the Appendix.
@@
Assumption 2: The training-test setting is always such that the model $\hat{f}$ converges at least with the parametric rate, i.e. $K\geq\overline{K}\;\Rightarrow\;\xi_{\mathrm{ml}}\geq\xi_{p}=\frac{1}{2}$ in (1).
@@ 3.4 Model Shapley values
Equations 3 & 4 do not account for the case when variables jointly contribute to model predictions, i.e. when they are dependent or interact. This situation can be addressed by using the more general Shapley-Taylor decomposition proposed by Agarwal et al. (2019): the discrete set derivative of model $\hat{f}$ at point $x_{i}$ with respect to the set $x^{\prime}$ conditioned on $x^{\prime\prime}$ is defined as
@@
$\phi_{h}\big{(}\hat{f},x_{i}\,|\,x^{\prime}\big{)}\,=\,\begin{cases}\delta_{x^{\prime}}\hat{f}(x_{i}\,|\,\emptyset)&\quad\text{if}\quad|x^{\prime}|<h\,,\\ \frac{h}{n}\sum_{x^{\prime\prime}\subseteq\mathcal{C}(x)\setminus x^{\prime}}\frac{\delta_{x^{\prime}}\hat{f}(x_{i}\,|\,x^{\prime\prime})}{\binom{n-1}{|x^{\prime\prime}|}}&\quad\text{if}\quad|x^{\prime}|=h\,.\end{cases}$
@@
with $\phi_{0}$ being the same intercept as in (3). The second sum generalizes summation over variables to that over Shapley-Taylor components. We suppress the expansion order $h$ in what follows if not explicitly needed. Statistical models usually do not allow for missing inputs, i.e. $|x^{\prime}|<n$. We address this by integrating out variables excluded from the model over a background $x_{\mathrm{bg}}$. The intercept $\phi_{0}$ is then the expected predicted value over $x_{\mathrm{bg}}$, i.e. $\phi_{0}=\mathbb{E}[\hat{f}(\emptyset)]$. This provides the reference value against which each Shapley component is measured. Thus, the choice of the background data is important for the interpretation of Shapley components, which we will discuss below. [^5] The formal properties of Shapley values, our computational approach, and a comparison of the Shapley-Taylor decomposition of model predictions and the the Taylor expansion of a analytical function are given in the Online Appendix.
@@ 4.1 Shapley estimation
$\bar{\phi}_{x^{\prime}}^{R}\,=\,\mathbb{E}_{\Omega}\big{[}\phi_{i,x^{\prime}}^{R}\big{]}=\frac{1}{m}\sum_{i=1}^{m}\,\phi_{i,x^{\prime}}^{R}\,.$ (8)
@@
(analytical continuity I) The Shapley decomposition $\Phi$ of a model $\hat{f}$ linear in parameters $\theta$, $\hat{f}(x)=x\theta$, is the model itself with $\Phi=\Phi_{1}$ (Eq. 3), and $\phi_{k}^{\mathrm{lin}}=(x_{k}-\bar{x}_{k})\hat{\theta}_{k}$ (Corollary 1 in Lundberg and Lee (2017)). The proof is given in the Appendix.
@@ 4.1.1 Application: Shapley estimation for heterogeneous treatment effects
$=$ $\bigg{[}\sum_{k=0}^{n}\phi_{i,k}^{t=1}+\sum_{k,l;k>0,k>l}\phi_{i,kl}^{t=1}\bigg{]}-\bigg{[}\sum_{k=0}^{n}\phi_{i,k}^{t=0}+\sum_{k,l;k>0,k>l}\phi_{i,kl}^{t=0}\bigg{]}$ (10)
@@
$=$ $\phi_{i,t}^{t=1}+\sum_{k;k\neq t}\phi_{i,tk}^{t=1}\;\equiv\;\phi_{t}+\sum_{k;k\neq t}\phi_{i,t*k}\,.\vspace*{-0.3cm}$ (12)
The second row (10) inserts the definition of the Shapley-Taylor decomposition where single and double indexed terms correspond to the main effects ($|x^{\prime}|=1$) and pairwise interactions ($|x^{\prime}|=2$), respectively. Any term not involving the treatment $t$ cancels out in the third row, including the intercept $\phi_{0}$.
Going to the forth row, we set the background data against which variable coalitions are evaluated to $x_{\mathrm{bg}}=(t=0,z)\equiv x_{\mathrm{bg}}^{0}$, i.e. the control group, or a representative subsample or summary of the untreated population. This has two consequences for for the treatment function. First, all terms in the second bracket of (11) vanish when going to (12) due to the missingness property of Shapley values. Second, the first terms $\phi_{t}$ in (12), which is constant across subjects, is an estimate of the average treatment effect (ATE).
@@ 4.2 Shapley estimation uncertainty
     $\cdot$ split $x$ into $x_{\mathrm{train}}^{s}$ and $x_{\mathrm{test}}^{s}$ using cross-fitting
@@
        $\cdot$ draw bootstrap sample $x_{b}^{s}$ from $x_{\mathrm{train}}^{s}$
@@
        $\cdot$ perform Shapley decomposition $\Phi^{b}(x_{\mathrm{test}}^{s};\hat{f}_{b})$
@@
$\sqrt{m}\,\bar{\phi}_{x^{\prime}}\;\;\underset{\mbox{\scriptsize$m\rightarrow\infty$}}{\text{\Large$\longrightarrow$}}\;\;\mathcal{N}\big{(}\bar{\phi}_{x^{\prime}}^{\star},\sigma_{x^{\prime}}^{\star}\big{)}\,.$
@@
(Sample split confidence intervals) Let $[\phi^{B}_{\mathrm{low}},\phi^{B}_{\mathrm{high}}]$ and $[\phi^{R}_{\mathrm{low}},\phi^{R}_{\mathrm{high}}]$ be confidence intervals at some level $\gamma$ from independent estimations, such as those resulting from bootstrap estimation (Algorithm 1) and sample split variation, respectively. Both intervals are median-centered at zero without loss of generality, then the joint confidence interval at level $\gamma$ is bounded as
$[\phi^{\mathrm{joint}}_{\mathrm{low}},\phi^{\mathrm{joint}}_{\mathrm{high}}]\;\subseteq\;[\phi^{B}_{\mathrm{low}}+\phi^{R}_{\mathrm{low}},\phi^{B}_{\mathrm{high}}+\phi^{R}_{\mathrm{high}}]\,.$ (14)
@@
The right-hand side of Equation 14 provides a conservative bound assuming that large variations in the estimation of $\phi^{B}$ coincides with large variations of $\phi^{R}$. However, when equating these quantities with sampling and sample split estimates, respectively, we often observe that the latter has considerably smaller variation than the former, i.e. $\phi^{R}_{\mathrm{high}}-\phi^{R}_{\mathrm{low}}\ll\phi^{B}_{\mathrm{high}}-\phi^{B}_{\mathrm{low}}$, such that right-hand side of (14) provides a practical adjustment. We will see in the empirical case study that such adjustments accounting for split uncertainty are mostly small, especially for aggregate measures where variation from the observation level cancels out.